# Benjamin S. Manning MIT John J. Horton

### v2

@@ 4 External validity in pre-committed novel settings
The setup is similar to that of Allcott (2015) and Hotz et al. (2005), where treatment effects from various “sites” are used to evaluate the external validity of a given intervention at the population level. However, they assume a common underlying intervention—analogous to a single setting in our framework—across all sites. And when there are heterogeneous interventions, only special instances with strong additional assumptions allow for appropriate inference. The following requires no such assumptions.
@@
The first two assumptions are basically identical to the assumptions of unconfounded location and random assignment in Hotz et al.. The first part of Assumption 3 is similar to the covariate overlap assumption in causal inference; without it, some observed responses would have $\log 0$. The second portion of Assumption 3 is a standard finite second-moment condition.
@@
The remainder of this section is devoted to implementing the above framework on a pre-committed family of strategic games. This set comprises 883,320 novel and unique permutations of Arad and Rubinstein’s money request game. We randomly sample 1,500 of these games for human subjects and AI subjects to play in a preregistered experiment. We use this data to estimate the relative capacity of different AI subjects to predict human responses at scale. In particular, we return to the strategic sample of optimized level-$k$ AI subjects presented in Table 2 from Section 3. We compare these agents’ ability to predict human responses across the 1,500 games to the baseline AI. We also calculate the unique symmetric Nash equilibria mechanically produced by a slightly modified version of the procedure in Harsanyi and Selten (1988) for each game and compare these theoretical predictions to the AI subjects. Because the games (and human subjects) are randomly sampled from the population according to a known distribution, confidence intervals over the comparisons are valid over the 883,320 games.
@@ 11. Player gets bonus if both players select the lower bound
Notes: This table shows the possible values for the six parameters of the pre-committed family of games. The naive Cartesian product of all parameter values yields $20\times 16\times 4\times 20\times 6\times 11=1{,}689{,}600$ combinations. However, seven of the eleven bonus rules do not use the gap parameter; for these rules, varying the gap value produces mechanically identical games. Collapsing such duplicates leaves 883,320 unique games in final population $X$.
@@
4.2 Eliciting AI subject responses
We generate AI responses for each of the 1,500 games in the set $S$ using two distinct samples of AI subjects. As a baseline sample, we prompt Gpt-4o at temperature 1 to independently play each game 100 times, without providing any additional instructions. For the optimized strategic sample, we use the same 10 prompts from Table 2, which were optimized using human experimental data from Arad and Rubinstein. To generate this strategic sample, we proportionally scale the optimized persona weights to create a total of 100 AI subjects, each of which plays each game exactly once using Gpt-4o (the “strategic” sample of AI subjects hereinafter). [^21] A very small fraction (less than 0.1%) of the AI subject responses were invalid due to stochasticity inherent to the LLM at temperature 1. Following our preregistered analysis plan, we discard these invalid responses without resampling.
This procedure produces an empirical distribution for both the strategic level-$k$ sample $\hat{P}_{\bm{\theta}^{*}}$ and the baseline AI $\hat{P}_{0}$ for every game $s\in S$. These samples correspond exactly to those described in Section 3, differing only in the number of agents—here, each distribution is generated with 100 agents rather than 1,000. To be clear, these agents were constructed months before they played any of the 1,500 games. In total, the elicitation procedure produces approximately 300,000 individual AI subject responses.
@@ 4.3 Harsanyi-Selten Nash equilibria as a benchmark
Another seemingly attractive alternative would be to follow the approaches of Fudenberg and Liang (2019) or Hirasawa et al. (2022), who train a bespoke supervised machine-learning model on past game data to predict responses. But, this method suffers from the same problem as the hierarchical models: it also relies heavily on having access to a large and representative dataset, which we currently do not possess.
@@
To implement this, we first calculate all Nash equilibria for the 1,500 games in set $S$ using open-source software (Savani and Turocy, 2025). We then apply the Harsanyi-Selten procedure to each game’s set of equilibria, producing a single equilibrium distribution $\hat{P}_{Nash}$ per game. This produced sets of equilibria for 1,487 games.222222 In fewer than 1% of the games, degeneracy issues prevented the code from converging. Following our preregistration, we discard these games in our analysis when comparing the equilibria to the strategic AI subjects. Of these 1,487 games, 467 have unique symmetric equilibria. The selection procedure was unnecessary in these cases. Among the remaining games with multiple symmetric equilibria, the procedure selects payoff-dominant equilibria—those Pareto superior to all alternatives—in 328 games, and risk-dominant equilibria in 1,026 games. Finally, 5959% of the equilibria selected by the Harsanyi-Selten procedure involve pure strategies, with the remainder employing mixed strategies.
@@ 4.4 Eliciting human responses
We collected human data from a sample of 4,500 Prolific workers using a custom online survey platform, which allows us to generate any game programmatically (Horton and Horton, 2024). This human data supplies $y_{s}$ for each game $s\in S$, with which we can then estimate the relative predictive power of the different AI models and equilibria. The entire experimental design, all AI subject responses, and the statistical analysis in this section were preregistered before collecting the human subjects’ data. Each prolific worker was randomly assigned one of the 1,500 sampled games such that each game had approximately three human players.
@@ 4.5 Estimation
We estimate the relative predictive power of the different AI samples and the theoretical benchmark in three steps. These are: (i) construct smoothed predictive distributions for every model in every sampled game; (ii) evaluate the strategic sample of AI subjects compared to each other model with per-game log-likelihoods and their paired differences; (iii) attach sampling-error bounds that are externally valid for the full population of nearly one million games. We address these steps in turn.
Many of the Harsanyi-Selten equilibria—mainly those pure strategies—and to a lesser extent the samples of AI subject place zero probability on strategies that humans sometimes take. For these models, the product of all likelihoods would be zero, making the log-likelihood -$\infty$ and violating Assumption 3. Dropping these games would heavily bias the results away from any pure strategy equilibria or the samples of AI subject where the agents mostly pick a single action—even when most people do select that action.
@@
For each game $s$, we calculate Equation 3 four times. In all four cases, $\theta^{\prime}$—the numerator of the log-likelihood ratio $\hat{\Lambda}_{s}$—is the sample of strategic AI subjects ($\hat{P}_{\bm{\theta}^{*}}$). The denominator $\theta^{\prime\prime}$ is one of four reference distributions: i) the baseline AI ($\hat{P}_{0}$), ii) the Harsanyi-Selten Nash equilibria ($\hat{P}_{Nash}$), iii) a uniform distribution over all possible strategies ($\hat{P}_{Unif}$), or iv) a randomly selected pure strategy distribution ($\hat{P}_{Pure}$). To be clear, this means all comparisons are made with respect to the strategic sample. We then take the average across the games to estimate $\bar{\Lambda}_{S}$ from Equation 4 for each comparison.
@@
We report bootstrapped confidence intervals for the four $\Lambda_{S}$ values and provide robustness checks with Wilcoxon and random-sign permutation tests. We also report the proportion of games for which the strategic AI subject is the best predictor—i.e. $\sum_{s\in S}\mathbf{1}\{\hat{\Lambda}_{s}>0\}/|S|$—with its exact Clopper-Pearson 95% interval. Such intervals are valid following a nearly identical argument leading to Proposition 1.
@@ 4.6 Results
Figure 6: Predictive power of Strategic AI subjects compared to other models ($\varepsilon=0.2$)
@@
Notes: The top panel shows the distribution of the log-likelihood ratios for each comparison. The vertical black line indicates the mean, which is also provided in the upper right corner of each panel. The respective bootstrapped standard errors are in parentheses. Green indicates ratios greater than zero, where the optimized AI has more predictive power, and red is the converse. The bottom panel shows the proportion of games for which the sample of strategic AI subjects is the best predictor. Standard errors are 95% Clopper-Pearson intervals.
@@
With $\varepsilon=0.2$, the strategic sample of AI subject is, on average, significantly more predictive than any of the other reference distributions ($p<0.001$ for all comparisons). These differences are substantial. Starting with the leftmost panel, across all human observations in the dataset, the strategic sample of AI subjects achieves an average per-observation likelihood ratio of $e^{1.23}=3.41$ in favor of the model, relative to the baseline AI. That is, the likelihood of the observed human data under the strategic AI subject model is, on average, $3.41$ times larger per observation than under the baseline. Moving right, the corresponding average likelihood ratios are $e^{0.89}=2.44$ against the Harsanyi–Selten-selected equilibria—this predictive edge is extant for both pure and mixed strategy equilibria (Table D5)—and $e^{0.59}=1.81$ against the uniform reference distribution. In the rightmost panel, the advantage over the random pure-strategy benchmark is largest, with an average per-observation likelihood ratio of $e^{4.15}=63.47$. These results are robust to Wilcoxon and random-sign permutation tests (Table D5).
@@
The bottom panel of Figure 6 shows the proportion of games for which the sample of strategic AI subjects is the best predictor—i.e. $\sum_{s\in S}\mathbf{1}\{\hat{\Lambda}_{s}>0\}/|S|$. The results are consistent with the top panel. The theoretically-motivated sample of strategic AI subjects better predicts the human responses in more games than any other model. This proportion is large and significant for the baseline ($0.715$). It is smaller, although still greater than 50%, for the Harsanyi-Selten Nash equilibria ($0.622$).
@@
Tables D3, D4, and D6 provide the same statistical analyses for $\varepsilon\in\{.05,0.1,0.3\}$, respectively. The above results are robust to these additional sensitivity checks. $\bar{\Lambda}_{S}>0$ for all comparisons, and the proportions of games for which the optimized AI subject is the best predictor are all greater than 50%. Tables D8, D9, D10, and D11 further show that the relative predictive power of the models is not particularly sensitive to the bonus or points rule. The strategic AI subjects are almost universally superior.
@@
Importantly, the strategic AI subjects also demonstrate impressive predictive power in the absolute. Without any smoothing, 24% of human respondents selected the strategy for which the optimized AI subject assigns the most density. Given that the number of possible strategies per game varied evenly between 5 and 20, this is notable. Furthermore, 53% of human respondents selected one of the top three strategies for which the strategic AI subject assigns the most density. And maybe most surprisingly, for 86% of games, all human respondents selected a strategy within the support of the strategic AI subject.
@@ 5 Conclusion
The great promise of AI subjects lies in their potential to accurately predict human behavior in novel settings. Realizing this capability could transform social science research and public policy. It could provide the social science equivalent of a lab bench in the physical sciences: an accurate, scalable playground to test ideas before large-scale and expensive implementation with humans. [^23] This could be even more powerful given the often observed researcher inability to predict results of their own experiments (DellaVigna et al., 2019; Milkman et al., 2021; Duckworth et al., 2025; Gandhi et al., 2024, 2023). Yet, as with the current state-of-the-art foundation models, AI subjects are not yet reliable enough to be used in this way out of the box.
@@
In this paper, we explored an approach to address this shortcoming. Our approach relies on two key principles: (i) grounding candidate AI subjects in theories expected to drive human behavior in the target setting, and (ii) optimizing and then validating AI subjects in distinct but related settings presumed to share underlying behavioral mechanisms. Without theoretical grounding, optimized prompts may fail to meaningfully improve even in-sample predictions. Without validation across distinct but related datasets, optimized prompts are prone to overfit a particular data-generating process. Just as economists carefully extend established theories to novel policy contexts—relying on accumulated empirical validation rather than absolute certainty—optimizing theoretically-grounded AI subjects to match samples of human data, and then validating them in distinct but related settings, provides a principled foundation for predicting humans in new settings.
@@
The improvements in predictive power yielded by this approach are substantial. In four novel and preregistered strategic games derived from Arad and Rubinstein’s 11-20 money request game, theoretically motivated AI subjects, optimized and validated through our methodology, reduced prediction errors by approximately 53-73% compared to baseline AI predictions. Remarkably, these theoretically-grounded agents predicted the results in some of the games better than the original human data from Arad and Rubinstein. Importantly, our results are not confined to games involving strategic reasoning. In Appendix A, we apply the same procedure to the allocation games from Charness and Rabin, and using data from a preregistered experiment with entirely novel human responses, find substantial improvements.
@@
Although this approach provides no statistical guarantees for arbitrary novel settings—indeed, no procedure can guarantee predictive power in entirely novel domains without a fully specified and correct causal model—we demonstrated that we can make externally valid inferences within a pre-committed family of settings. Using novel data from 4,249 participants playing 1,490 games randomly sampled from a heterogeneous population of 883,320 strategic games, we found that the strategic level-$k$ agents generalized effectively across this broad domain. In 86% of games, all human subjects chose actions within the support of the optimized AI subject responses. These general agents substantially outperformed the baseline AI off-the-shelf, the Harsanyi-Selten theoretical predictions, and several other benchmarks. The results were robust to several alternative specifications. Because the games were randomly sampled under the assumptions stated in Section 4.5, these results are externally valid for the entire population of 883,320 games. While we can only guarantee validity within this specific population, the strong performance of theoretically motivated AI subjects across such a diverse set of strategic games suggests that similar approaches would likely generalize to other game-theoretic settings.
@@
More generally, the results in this paper linking theoretically motivated AI subjects to robust generalizability in novel settings, and atheoretical AI subjects to failure, are notable for reasons beyond prediction (Hofman et al., 2021). They suggest that the underlying LLM has correctly learned the relevant relationships between the AI subjects and human responses to the given setting. This is even more notable given that it is highly unlikely that such a mapping was explicitly specified during training. Such a finding aligns with recent evidence that LLMs form rich internal representations of their human-generated training corpus rather than merely memorizing it (Lindsey et al., 2025; Ameisen et al., 2025). If true for LLMs and human behavior more generally, our prompt-alignment-generalizability exercises may offer more than improved predictive power. If an LLM armed with a particular theoretically-motivated prompt matches human data particularly well across a wide range of related settings, it might be evidence that the theory has a lot of explanatory power for the underlying human sample. Building on a young social scientific literature (Si et al., 2024; Ludwig and Mullainathan, 2024; Mullainathan and Rambachan, 2024; Batista and Ross, 2024; Enke and Shubatt, 2023; Peterson et al., 2021; Movva et al., 2025), this could, in turn, provide researchers with robust machine-learning methods to rapidly and efficiently inform promising new hypotheses.

### v3

@@ 4 External validity in pre-committed novel settings
The setup is related to that of Hotz et al. (2005), Allcott (2015), and Dehejia et al. (2019), where treatment effects from various “sites” are used to evaluate the external validity of a given intervention at the population level. However, they assume a common underlying intervention—analogous to a single setting in our framework—across all sites. And when there are heterogeneous interventions, only special instances with strong additional assumptions allow for appropriate inference. The following, which is very similar to the theoretical framework in Andrews et al. (2025), requires no such assumptions.
@@
The first two assumptions are basically identical to the assumptions of unconfounded location and random assignment in Hotz et al.. They are also very similar in spirit to Assumption 1 in Andrews et al.. The first part of Assumption 3 is similar to the covariate overlap assumption in causal inference; without it, some observed responses would have $\log 0$. The second portion of Assumption 3 is a standard finite second-moment condition.
@@
The remainder of this section is devoted to implementing the above framework on a pre-committed family of strategic games. This set comprises 883,320 novel and unique permutations of Arad and Rubinstein’s money request game. We randomly sample 1,500 of these games for human subjects and AI agents to play in a preregistered experiment. We use this data to estimate the relative capacity of different AI agents to predict human responses at scale. In particular, we return to the strategic sample of optimized level-$k$ AI agents presented in Table 2 from Section 3. We compare these agents’ ability to predict human responses across the 1,500 games to the baseline AI. We also calculate the unique symmetric Nash equilibria mechanically produced by a slightly modified version of the procedure in Harsanyi and Selten (1988) for each game and compare these theoretical predictions to the AI agents. Because the games (and human subjects) are randomly sampled from the population according to a known distribution, confidence intervals over the comparisons are valid over the 883,320 games.
@@ 11. Player gets bonus if both players select the lower bound
Notes: This table shows the possible values for the six parameters of the pre-committed family of games. The naive Cartesian product of all parameter values yields $20\times 16\times 4\times 20\times 6\times 11=1{,}689{,}600$ combinations. However, seven of the eleven bonus rules do not use the gap parameter; for these rules, varying the gap value produces mechanically identical games. Collapsing such duplicates leaves 883,320 unique games in the final population $X$.
@@
4.2 Eliciting AI agent responses
We generate AI responses for each of the 1,500 games in the set $S$ using two distinct samples of AI agents. As a baseline sample, we prompt Gpt-4o at temperature 1 to independently play each game 100 times, without providing any additional instructions. For the optimized strategic sample, we use the same 10 prompts from Table 2, which were optimized using human experimental data from Arad and Rubinstein. To generate this strategic sample, we proportionally scale the optimized persona weights to create a total of 100 AI agents, each of which plays each game exactly once using Gpt-4o (the “strategic” sample of AI agents hereinafter). [^21] A very small fraction (less than 0.1%) of the AI agent responses were invalid due to stochasticity inherent to the LLM at temperature 1. Following our preregistered analysis plan, we discard these invalid responses without resampling.
This procedure produces an empirical distribution for both the strategic level-$k$ sample $\hat{P}_{\bm{\theta}^{*}}$ and the baseline AI $\hat{P}_{0}$ for every game $s\in S$. These samples correspond exactly to those described in Section 3, differing only in the number of agents—here, each distribution is generated with 100 agents rather than 1,000. To be clear, these agents were constructed months before they played any of the 1,500 games. In total, the elicitation procedure produces approximately 300,000 individual AI agent responses.
@@ 4.3 Harsanyi-Selten Nash equilibria as a benchmark
Another seemingly attractive alternative would be to follow the approaches of Fudenberg and Liang (2019), Hirasawa et al. (2022), or Andrews et al. (2025), which involve training a bespoke supervised machine-learning model on past game data to predict responses. However, this method is problematic, if not infeasible, for two reasons. First, it suffers from the same issue as the hierarchical models: it relies heavily on having access to a large and representative dataset, which we currently do not possess. Second, as discussed in the motivating example from Section 2, traditional machine learning models (e.g., regression, decision trees, or generic neural networks) are not well-suited for predicting human behavior in novel settings because they cannot adapt to structural differences across settings. Even with substantial representative data, if a game had even one new parameter value (e.g., a new bonus rule or an upper bound above the highest from the training data), the model would need to be retrained from scratch.
@@
To implement this, we first calculate all Nash equilibria for the 1,500 games in set $S$ using open-source software (Savani and Turocy, 2025). We then apply the Harsanyi-Selten procedure to each game’s set of equilibria, producing a single equilibrium distribution $\hat{P}_{Nash}$ per game. This produced sets of equilibria for 1,487 games.222222 In fewer than 1% of the games, degeneracy issues prevented the code from converging. Following our preregistration, we discard these games in our analysis when comparing the equilibria to the strategic AI agents. Of these 1,487 games, 467 have unique symmetric equilibria. The selection procedure was unnecessary in these cases. Among the remaining games with multiple symmetric equilibria, the procedure selects payoff-dominant equilibria—those Pareto superior to all alternatives—in 328 games, and risk-dominant equilibria in 1,026 games. Finally, 5959% of the equilibria selected by the Harsanyi-Selten procedure involve pure strategies, with the remainder employing mixed strategies.
@@ 4.4 Eliciting human responses
We collected human data from a sample of 4,500 Prolific workers using a custom online survey platform, which allows us to generate any game programmatically (Horton and Horton, 2024). This human data supplies $y_{s}$ for each game $s\in S$, with which we can then estimate the relative predictive power of the different AI models and equilibria. The entire experimental design, all AI agent responses, and the statistical analysis in this section were preregistered before collecting the human subjects’ data. Each prolific worker was randomly assigned one of the 1,500 sampled games such that each game had approximately three human players.
@@ 4.5 Estimation
We estimate the relative predictive power of the different AI samples and the theoretical benchmark in three steps. These are: (i) construct smoothed predictive distributions for every model in every sampled game; (ii) evaluate the strategic sample of AI agents compared to each other model with per-game log-likelihoods and their paired differences; (iii) attach sampling-error bounds that are externally valid for the full population of nearly one million games. We address these steps in turn.
Many of the Harsanyi-Selten equilibria—mainly those pure strategies—and to a lesser extent the samples of AI agent place zero probability on strategies that humans sometimes take. For these models, the product of all likelihoods would be zero, making the log-likelihood -$\infty$ and violating Assumption 3. Dropping these games would heavily bias the results away from any pure strategy equilibria or the samples of AI agent where the agents mostly pick a single action—even when most people do select that action.
@@
For each game $s$, we calculate Equation 3 four times. In all four cases, $\theta^{\prime}$—the numerator of the log-likelihood ratio $\hat{\Lambda}_{s}$—is the sample of strategic AI agents ($\hat{P}_{\bm{\theta}^{*}}$). The denominator $\theta^{\prime\prime}$ is one of four reference distributions: i) the baseline AI ($\hat{P}_{0}$), ii) the Harsanyi-Selten Nash equilibria ($\hat{P}_{Nash}$), iii) a uniform distribution over all possible strategies ($\hat{P}_{Unif}$), or iv) a randomly selected pure strategy distribution ($\hat{P}_{Pure}$). To be clear, this means all comparisons are made with respect to the strategic sample. We then take the average across the games to estimate $\bar{\Lambda}_{S}$ from Equation 4 for each comparison.
@@
We report bootstrapped confidence intervals for the four $\Lambda_{S}$ values and provide robustness checks with Wilcoxon and random-sign permutation tests. We also report the proportion of games for which the strategic AI agent is the best predictor—i.e. $\sum_{s\in S}\mathbf{1}\{\hat{\Lambda}_{s}>0\}/|S|$—with its exact Clopper-Pearson 95% interval. Such intervals are valid following a nearly identical argument leading to Proposition 1.
@@ 4.6 Results
Figure 6: Predictive power of Strategic AI agents compared to other models ($\varepsilon=0.2$)
@@
Notes: The top panel shows the distribution of the log-likelihood ratios for each comparison. The vertical black line indicates the mean, which is also provided in the upper right corner of each panel. The respective bootstrapped standard errors are in parentheses. Green indicates ratios greater than zero, where the optimized AI has more predictive power, and red is the converse. The bottom panel shows the proportion of games for which the sample of strategic AI agents is the best predictor. Standard errors are 95% Clopper-Pearson intervals.
@@
With $\varepsilon=0.2$, the strategic sample of AI agent is, on average, significantly more predictive than any of the other reference distributions ($p<0.001$ for all comparisons). These differences are substantial. Starting with the leftmost panel, across all human observations in the dataset, the strategic sample of AI agents achieves an average per-observation likelihood ratio of $e^{1.23}=3.41$ in favor of the model, relative to the baseline AI. That is, the likelihood of the observed human data under the strategic AI agent model is, on average, $3.41$ times larger per observation than under the baseline. Moving right, the corresponding average likelihood ratios are $e^{0.89}=2.44$ against the Harsanyi–Selten-selected equilibria—this predictive edge is extant for both pure and mixed strategy equilibria (Table D5)—and $e^{0.59}=1.81$ against the uniform reference distribution. In the rightmost panel, the advantage over the random pure-strategy benchmark is largest, with an average per-observation likelihood ratio of $e^{4.15}=63.47$. These results are robust to Wilcoxon and random-sign permutation tests (Table D5).
@@
The bottom panel of Figure 6 shows the proportion of games for which the sample of strategic AI agents is the best predictor—i.e. $\sum_{s\in S}\mathbf{1}\{\hat{\Lambda}_{s}>0\}/|S|$. The results are consistent with the top panel. The theory-grounded sample of strategic AI agents better predicts the human responses in more games than any other model. This proportion is large and significant for the baseline ($0.715$). It is smaller, although still greater than 50%, for the Harsanyi-Selten Nash equilibria ($0.622$).
@@
Tables D3, D4, and D6 provide the same statistical analyses for $\varepsilon\in\{.05,0.1,0.3\}$, respectively. The above results are robust to these additional sensitivity checks. $\bar{\Lambda}_{S}>0$ for all comparisons, and the proportions of games for which the optimized AI agent is the best predictor are all greater than 50%. Tables D8, D9, D10, and D11 further show that the relative predictive power of the models is not particularly sensitive to the bonus or points rule. The strategic AI agents are almost universally superior across different slices of the parameter space in Table 3.
@@
Importantly, the strategic AI agents also demonstrate impressive predictive power in the absolute. Without any smoothing, 24% of human respondents selected the strategy for which the optimized AI agent assigns the most density. Given that the number of possible strategies per game varied evenly between 5 and 20, this is notable. Furthermore, 53% of human respondents selected one of the top three strategies for which the strategic AI agent assigns the most density. And maybe most surprisingly, for 86% of games, all human respondents selected a strategy within the support of the strategic AI agent.
@@ 5 Conclusion
The great promise of AI agents lies in their potential to accurately predict human behavior in novel settings. Realizing this capability could transform social science research and public policy. It could provide the social science equivalent of a lab bench in the physical sciences: an accurate, scalable playground to test ideas before large-scale and expensive implementation with humans. [^23] This could be even more powerful given the often observed researcher inability to predict results of their own experiments (DellaVigna et al., 2019; Milkman et al., 2021; Duckworth et al., 2025; Gandhi et al., 2024, 2023). Yet, as with the current state-of-the-art foundation models, AI agents are not yet reliable enough to be used in this way out of the box.
@@
In this paper, we explored an approach to address this shortcoming. Our approach relies on two key principles: (i) grounding candidate AI agents in theories expected to drive human behavior in the target setting, and (ii) optimizing and then validating AI agents in distinct but related settings presumed to share underlying behavioral mechanisms. Without theoretical grounding, optimized prompts may fail to meaningfully improve even in-sample predictions. Without validation across distinct but related datasets, optimized prompts are prone to overfit a particular data-generating process. Just as economists carefully extend established theories to novel policy contexts—relying on accumulated empirical validation rather than absolute certainty—optimizing theoretically-grounded AI agents to match samples of human data, and then validating them in distinct but related settings, provides a principled foundation for predicting humans in new settings.
@@
The improvements in predictive power yielded by this approach are substantial. In four novel and preregistered strategic games derived from Arad and Rubinstein’s 11-20 money request game, theoretically motivated AI agents, optimized and validated through our methodology, reduced prediction errors by approximately 53-73% compared to baseline AI predictions. Remarkably, these theoretically-grounded agents predicted the results in some of the games better than the original human data from Arad and Rubinstein. Importantly, our results are not confined to games involving strategic reasoning. In Appendix A, we apply the same procedure to the allocation games from Charness and Rabin, and using data from a preregistered experiment with entirely novel human responses, find substantial improvements.
@@
Although this approach provides no statistical guarantees for arbitrary novel settings—indeed, no procedure can guarantee predictive power in entirely novel domains without a fully specified and correct causal model—we demonstrated that we can make externally valid inferences within a pre-committed family of settings. Using novel data from 4,249 participants playing 1,490 games randomly sampled from a heterogeneous population of 883,320 strategic games, we found that the strategic level-$k$ agents generalized effectively across this broad domain. In 86% of games, all human subjects chose actions within the support of the optimized AI agent responses. These general agents substantially outperformed the baseline AI off-the-shelf, the Harsanyi-Selten theoretical predictions, and several other benchmarks. The results were robust to several alternative specifications. Because the games were randomly sampled under the assumptions stated in Section 4.5, these results are externally valid for the entire population of 883,320 games. While we can only guarantee validity within this specific population, the strong performance of theoretically motivated AI agents across such a diverse set of strategic games suggests that similar approaches would likely generalize to other game-theoretic settings.
@@
More generally, the results in this paper linking theoretically motivated AI agents to robust generalizability in novel settings, and atheoretical AI agents to failure, are notable for reasons beyond prediction (Hofman et al., 2021). They suggest that the underlying LLM has correctly learned the relevant relationships between the AI agents and human responses to the given setting. This is even more notable given that it is highly unlikely that such a mapping was explicitly specified during training. Such a finding aligns with recent evidence that LLMs form rich internal representations of their human-generated training corpus rather than merely memorizing it (Lindsey et al., 2025; Ameisen et al., 2025). If true for LLMs and human behavior more generally, our prompt-alignment-generalizability exercises may offer more than improved predictive power. If an LLM armed with a particular theory-grounded prompt matches human data particularly well across a wide range of related settings, it might be evidence that the theory has a lot of explanatory power for the underlying human sample. Building on a young social scientific literature (Si et al., 2024; Ludwig and Mullainathan, 2024; Mullainathan and Rambachan, 2024; Batista and Ross, 2024; Enke and Shubatt, 2023; Peterson et al., 2021; Movva et al., 2025; Hirasawa et al., 2022), this could, in turn, provide researchers with robust machine-learning methods to rapidly and efficiently inform promising new hypotheses.